On 1 July 2019, Weiland Pty Ltd owned several farming vehicles that had cost a total of $155 000. Accumulated depreciation on these vehicles to 1 July 2019 amounted to $73 000. On 30 September 2019, Weiland Pty Ltd acquired a new delivery vehicle and traded in one vehicle that had cost $32 000 and which had accumulated depreciation of $14 400 up to 1 July 2019. The full price of the new vehicle was $29 500 and the trade-in value of the old vehicle was agreed at $18 400. Ignore GST. On 31 December 2019, an additional delivery vehicle was purchased for $34 000 cash. Depreciation is calculated at the rate of 20% p.a. on the diminishing balance.

1.Prepare the appropriate journal entries for the year ended 30 June 2020
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Weiland Pty Ltd Workings
Vehicle sold on 30 September 2019
Particulars Amount $ Note
Cost as on 1st July 2019     32,000.00
Less: Accumulated Depreciation as on 1st July 2019     14,400.00 V
Book Value as on 1st July 2019     17,600.00 A
Depreciation rate 20% B
Depreciation for 2019       3,520.00 C=A*B
Number of months from 1st July to 30th September               3.00 D
Depreciation for 3 months          880.00 E=C/12*D
Book Value as on 30th September 2019     16,720.00 F=A-E
Trade in Value     18,400.00 G
Profit on exchange       1,680.00 H=G-F
Accumulated Depreciation as on 30th September 2019     15,280.00 U=V+E
Vehicle purchased on 30 Sept 2019
Particulars Amount $ Note
Cost of Machine     29,500.00 K
Less: Trade in Value     18,400.00
Cash paid     11,100.00 I
Depreciation rate 20% See B
Depreciation for 2019       5,900.00 J=K*B
Number of months from 30th September to 30th June 2020               9.00 L
Depreciation for 9 months       4,425.00 M=J/12*L
Vehicle purchased on 31st Dec 2019
Particulars Amount $ Note
Cost of Machine     34,000.00 N
Depreciation rate 20% See B
Depreciation for 2019       6,800.00 O=N*B
Number of months from 31st Dec to 30th June 2020               6.00 P
Depreciation for 6 months       3,400.00 Q=O/12*P
Depreciation on opening book value of Vehicle
Particulars Amount $ Note
Cost as on 1st July 2019 155,000.00
Less: Accumulated Depreciation as on 1st July 2019     73,000.00
Book Value as on 1st July 2019     82,000.00 R
Less: Book value of vehicle sold on 30 September 2019     17,600.00 See A
Depreciable value     64,400.00 S
Depreciation rate 20% See B
Depreciation for 2019     12,880.00 T=S*B
Journal Entries
Date Account Debit $ Credit $
Sep. 30, 2019 Depreciation        880.00
Sep. 30, 2019 Accumulated Depreciation        880.00
(See calculation in Note E.)
Sep. 30, 2019 Vehicle purchased 29,500.00 (See Note K.)
Sep. 30, 2019 Accumulated Depreciation 15,280.00 (See calculation in Note U.)
Sep. 30, 2019 Vehicle sold 32,000.00 (See Note G.)
Sep. 30, 2019 Profit & Loss     1,680.00 (See calculation in Note H.)
Sep. 30, 2019 Cash 11,100.00 (See calculation in Note I.)
June. 30, 2020 Depreciation 20,705.00 (See calculation in Note W.)
June. 30, 2020 Accumulated Depreciation 20,705.00